OpenStudy (anonymous):

Factor the following Polynomials. 25z^2-1 4c^2-121 9a^2-100 I'm not sure how to do this at all. I probably have an idea but I'm bot sure what it means or how to do it.

OpenStudy (anonymous):

Help please!!

OpenStudy (anonymous):

use quadratic formula

OpenStudy (raden):

all questions above can be factored by using difference square formula : a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

so like #1 (25-1)(25+1)??

OpenStudy (anonymous):

(5z−1)(5z+1)

OpenStudy (anonymous):

did you square root 25?

OpenStudy (anonymous):

yes.

OpenStudy (raden):

25z^2-1 = (5z)^2 - 1^2 = (5z - 1)(5z + 1)

OpenStudy (anonymous):

ok sweet i think I got it. so would the 4c^2-121 be (2c+11)(2c-11)

OpenStudy (anonymous):

Yes you did it right.

OpenStudy (anonymous):

2A^4-18a^2?

OpenStudy (raden):

2A^4-18a^2 = 2 (A^4 - 9a^2) = 2 ( (A^2)^2 - (3a)^2) ) = 2 (A^2 + 3a)(A^2 - 3a)

OpenStudy (anonymous):

2a^2(a-3)(a+3)

OpenStudy (raden):

i assumed that "A" different with "a"

OpenStudy (raden):

so, just 2 (A^2 + 3a)(A^2 - 3a)

OpenStudy (anonymous):

Ohh lol i didn't notice. Your totally right. (:

OpenStudy (raden):

:*